OpenStudy (anonymous):

Find y'' by implicit differentiation 5x^2+y^2=6

OpenStudy (pixiedust1):

Welcome to OpenStudy! :)

OpenStudy (anonymous):

Thanks :)

OpenStudy (mathmale):

1. apply the derivative operator (d/dx) to each of the three terms in the original equation. Note that we regard y as a function of x, and therefore must use the chain as well as the power rule in finding the derivative with respect to x of y^2. 2. Solve the resulting equation for the first derivative, y '(x). 3. Apply the derivative operator to y '(x), 4. Simplify the result. Good luck! I'm loggiing off, but hope this helps you get started!

OpenStudy (anonymous):

no wait

OpenStudy (mathmale):

Label your final answer as y ''(x).

OpenStudy (anonymous):

i got the first derivative -5x/y

OpenStudy (anonymous):

could you help me with the second derivative?

OpenStudy (amistre64):

you found y', but its easier (for me) to play with y'' format if you retain the implicit sturcture 5x^2 + y^2 = 6 5x + y y' = 0 derive again

OpenStudy (anonymous):

\[5x^2+y^2=6,5*2x+2yy'=0,5x+yy'=0\]

OpenStudy (mathmale):

Be certain to label that first derivative! I agree that it may be faster to find the second derivative by leaving the result of the first differentiation in implicit form. Note that this requires that you use the PRODUCT rule, as well as the CHAIN rule.

OpenStudy (anonymous):

i cant seem to get it right.. probably making some mistake in between

OpenStudy (amistre64):

type out your steps :)

OpenStudy (amistre64):

\[\Large \underbrace{5x}_{power~rule} + \underbrace{y y'}_{product/chain} = \underbrace{0}_{constant}\]

OpenStudy (anonymous):

is y'=-5x/y ?

OpenStudy (amistre64):

yes, y' = -5x/y is good to keep in mind

OpenStudy (anonymous):

5x+y.(-5x/y)=0 ?

OpenStudy (amistre64):

or quotient rule:\[\frac tb\to\frac{bt'-b't}{b^2}\]

OpenStudy (amistre64):

not sure if we should subsititute y' in the middle of it, i tend to do it at the end. 5x + y y' = 0 5 + [y' y' + y y''] = 0

OpenStudy (anonymous):

so 5 + [(-5x/y)(-5x/y)+(y)(y'')=0 ? but i don't know what y'' is

OpenStudy (amistre64):

y'' is some unknown that we algebra, in the same way you algebraed y'

OpenStudy (amistre64):

5 + y'y' + y y'' = 0 5 + (y')^2 + y y'' = 0 y y'' = -(5+(y')^2) y'' = -(5+(y')^2)/y

OpenStudy (amistre64):

how much you want to simplify it is up to you :)

OpenStudy (anonymous):

thanks.. i'm trying to get the correct answer

OpenStudy (amistre64):

\[y'=\frac{-5x}{y}~~\to ~~y''=\frac{-5y+5xy'}{y^2}\] the correct answer depends on how far you want to simplify it

OpenStudy (anonymous):

what should i substitute with y?

OpenStudy (amistre64):

y is just y ... just as x is just x. they tend to be place holders for some (x,y) point from the original function

OpenStudy (amistre64):

in other words, an implicit derivative maintains the possibility that any derivative is a function of x and y

OpenStudy (amistre64):

just as y' = -5x/y is a function in x and y, for some known (x,y) point that satisfies the original setup

OpenStudy (anonymous):

-30/y^3 ??

OpenStudy (amistre64):

maybe .. lets see. \[y'' = -\frac{5+(y')^2}{y}\] \[y'' = -\frac{5+(-5\frac xy)^2}{y}\] \[y'' = -\frac{5+25\frac {x^2}{y^2}}{y}\] \[y'' = -\frac{5y^2+25x^2}{y~y^2}\] \[y'' = -5\frac{y^2+5x^2}{y^3}\]

OpenStudy (amistre64):

or \[y'=\frac{-5x}{y}~~\to ~~y''=\frac{-5y+5xy'}{y^2}\] \[y''=\frac{-5y+5x(-5\frac xy)}{y^2}\] \[y''=\frac{-5y-25\frac{x^2}{y}}{y^2}\] \[y''=\frac{-5y^2-25x^2}{y^2~y}\] \[y''=-5\frac{y^2+5x^2}{y^3}\]

OpenStudy (anonymous):

the correct answer is -30/y^3, is there any way u can get that answer from the value of y'' above?

OpenStudy (anonymous):

??

OpenStudy (amistre64):

where are you getting that 'correct' answer from?

OpenStudy (amistre64):

and no, i have shown 2 ways to approach the correct answer and they both match. so this -30/y^3 seems suspect to me

OpenStudy (anonymous):

i entered the answer according to your explanation for y'' but it is wrong, i'm doing it online and have used all 3 of my submissions so it gives the answer after no more tries are left

OpenStudy (anonymous):

i think there is some other way

OpenStudy (amistre64):

can you post the entire question, i think you are mixing information here or leaving something out.

OpenStudy (amistre64):

even the wolf agrees with my assessment http://www.wolframalpha.com/input/?i=%285+%28x+%28-5x%2Fy%29-y%29%29%2Fy%5E2

OpenStudy (anonymous):

i mean i know i'm not gonna get the points back for that question but i still want to know the right one

OpenStudy (anonymous):

i know i totally understand your explanation

OpenStudy (amistre64):

take a screen shot, and attach it please

OpenStudy (anonymous):

but i don't know what the question wants

OpenStudy (anonymous):

i'm working on the second question now

OpenStudy (amistre64):

screen shot that one then, lets take the middle man out of the setup :)

OpenStudy (anonymous):

i dont know how to :/

OpenStudy (anonymous):

but seriously the answer is -30/y^3 and questions is asking to find y'' by implicit differentiation

OpenStudy (amistre64):

on your keyboard is a button that youve never used, PrtScrn, its aroung hte scroll lock button .... press it. open a paint program and paste it in. save the file and attach it

OpenStudy (amistre64):

y' = -5x/y is correct, so the y'' is gonna have to have some sort of x in it, so i have to wonder about the content of the question, or the validity of the grading program :)

OpenStudy (anonymous):

okay i know this questions is confusing me a lot

OpenStudy (amistre64):

your input was off, even tho the 'answer' is wrong by a long shot lol \[\large y''=-5~\frac{y^2+5x^2}{y^3}\] or \[\large y''=\frac{-5y^2-25x^2}{y^3}\] NOT \[\large y''=-5y-\frac{25x^2}{y^2}\]

OpenStudy (amistre64):

but yeah ... thats messed up regardless lol

OpenStudy (anonymous):

i know i did make a mistake but that was before you showed me the right way lol

OpenStudy (anonymous):

i know right

OpenStudy (amistre64):

show it to your teacher .... if the 1pt is that important ;)

OpenStudy (anonymous):

so what do i do ???? i have more questions like this

OpenStudy (anonymous):

i have like 8 questions on this topic and that 1 point constitutes to about 13% of the whole'

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

brb in 2 mins... dont go!!

OpenStudy (amistre64):

y' = -5x/y is good or in product form y' = -5x y^(-1) y'' = -5y^(-1) + 5xy' y^(-2) substitute in y' y'' = -5y^(-1) + 5x(-5x/y) y^(-2) y'' = -5/y -25x^2/y^3 y'' = [-5y^2 - 25x^2 ]/y^3

OpenStudy (amistre64):

ill stay here, gonna do some desk work

OpenStudy (amistre64):

now that you know how to screenshot, feel free to attach another one

OpenStudy (anonymous):

back.. and yeah thanks so much for that :)

OpenStudy (anonymous):

ok so here's another one its super similar to the previous one lol

OpenStudy (amistre64):

then lets make a general rule :) ax^2 + by^2 = c 2ax x' + 2b y y' = 0 ; divide off the 2s, and x' = dx/dx = 1 ax + by y' = 0, y' = -ax/by continue with y'' .... ax + by y' = 0 ax' + by' y' + by y'' = 0; x'=1 again a + b(y')^2 + by y'' = 0 by y'' = -a - b(y')^2 y'' = [-a - b(y')^2 ] / by plug in y' y'' = [-a - b(-ax/by)^2 ] / by y'' = [-a - b(a^2x^2/b^2y^2) ] / by y'' = [-a - a^2x^2/by^2 ] / by y'' = [-aby^2 - a^2x^2 ] / [b^2 y^3]

OpenStudy (amistre64):

so, a=4, b=1, c=3 y'' = [-4y^2 - 16x^2 ] / [y^3] or as last time, a=5, b=1, c= whocares its not in the end y'' = [-5y^2 - 25x^2 ] / [y^3]

OpenStudy (amistre64):

now, according to the correctness of the last one, the solution HAS TO be -20/y^3 lol

OpenStudy (anonymous):

let me solve it myself.. give me minute

OpenStudy (amistre64):

k :)

OpenStudy (anonymous):

lol how did u come up with the solution of -20/y^3?

OpenStudy (amistre64):

well, since -5-25 = -30 i just assume their ignorance and let -4-16 = -20

OpenStudy (anonymous):

that isn't correct though lol

OpenStudy (amistre64):

the correct answer is .... \[\large y'' = \frac{-4y^2 - 16x^2 }{y^3}\]

OpenStudy (amistre64):

what they want the correct answer to be is anyones guess :)

OpenStudy (amistre64):

theres a watch it option, click it to see what happens

OpenStudy (anonymous):

i got it wrong again... it was -12/y^3

OpenStudy (anonymous):

the 'watch it' option just takes me to a video where a guy is solving for y' but not y''

OpenStudy (amistre64):

x^2 y^2 ---- + --- = 1 3/4 3/1  at the point 0,sqrt(3) y' = -4(0)/sqrt(3) = 0 as expected y'' = (-4(0)-16(3))/(3sqrt(3)) y'' = -16/sqrt(3), which is the rate at which the slope is changing